OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x-2y = 11

sam (.sam.):

5x-2y = 11 -2y=11-5x y=-11/2 + 5x/2 gradient or slope or m, is 5/2

OpenStudy (anonymous):

parallel lines have the same gradient : -theory